A car is to make a term with a radius of curvature of 60 m at a speed of 27 m/s. At what angles should the road be banked for the car to make the turn?

1
Expert's answer
2021-11-18T12:22:06-0500

The angle is


"\theta=\arctan\frac{v^2}{rg}=51°"